# On Sobolev and Besov spaces with hybrid regularity

### v1

@@ 3.1. Sobolev Spaces
we get a primal basis which is normalised in $L^{p}(\square)$, and a dual basis which is normalised in $L^{p^{\prime}}(\square)$, where $\frac{1}{p}+\frac{1}{p^{\prime}}=1$. With the above notation at hand, every function $u\in L^{p}(\square)$ admits unique expansions
[…]
@@
$\|u\|_{H^{t}(\square)}^{2}\sim\sum_{\boldsymbol{\lambda}\in\boldsymbol{\nabla}}2^{2t|\boldsymbol{\lambda}|_{\infty}}\big{|}\langle u,\psi_{\boldsymbol{\lambda}}\rangle\big{|}^{2}\sim\sum_{\mu\in\hexagon}2^{2s|\mu|}\big{\langle}u,\theta_{\mu}\rangle\big{|}^{2},\quad-\gamma<t<\tilde{\gamma},$ (13)
cf. [1, 6, 16, 19, 25, 26]. In particular, a function $u$ is contained in an isotropic Sovolev space if the coefficients of its (tensor-product) wavelet expansion decay sufficiently fast. Note that we have slightly abused the notation here, as the primal and dual wavelets may characterise function spaces with complementary boundary conditions.
@@
By using the norm equivalences (12) and (13), we see that we can also characterise the spaces $\mathfrak{H}^{q,s}(\square)$ in terms of wavelet coefficients.
@@
$\|u\|_{H^{q+s}([0,1])\otimes H^{s}([0,1])}^{2}$ $\sim\sum_{\boldsymbol{\lambda}\in\boldsymbol{\nabla}}2^{2(q+s)|\lambda_{\mathrm{x}}|+2s|\lambda_{\mathrm{y}}|}|u_{\boldsymbol{\lambda}}|^{2},$
$\|u\|_{H^{s}([0,1])\otimes H^{q+s}([0,1])}^{2}$ $\sim\sum_{\boldsymbol{\lambda}\in\boldsymbol{\nabla}}2^{2s|\lambda_{\mathrm{x}}|+2(q+s)|\lambda_{\mathrm{y}}|}|u_{\boldsymbol{\lambda}}|^{2},$
provided that $-\tilde{\gamma}<s,q+s<\gamma$. Therfore, as $q+s\geq s$, there holds
$2^{2(q+s)|\lambda_{\mathrm{x}}|+2s|\lambda_{\mathrm{y}}|}+2^{2s|\lambda_{\mathrm{x}}|+2(q+s)|\lambda_{\mathrm{y}}|}$ $\sim 2^{2(q+s)|\boldsymbol{\lambda}|_{\infty}+2s\min\{|\lambda_{\mathrm{x}}|,|\lambda_{\mathrm{y}}|\}}$
@@
where we have used (16) for $-q\geq 0$.
@@
Note that we have again used (16) for $-q\geq 0$. This shows (15).
@@
For a smaller range of parameters, the statement of Theorem 3.3 has already been established in [15]. Moreover, we note that the one-sided upper bound in (15) can be extended for $-\tilde{d}<s,q+s<\gamma$, whereas the one-sided lower bound can be extended to $-\tilde{\gamma}<s,q+s<d$. Similarly, we can also extend the upper and lower bounds in (16) up to $-d$ and $\tilde{d}$, respectively.
We note that there holds $\theta_{\mu},\psi_{\boldsymbol{\lambda}}\in\mathfrak{H}^{0,\gamma}(\square)\subseteq\mathfrak{H}^{\gamma,0}(\square)=H^{\gamma}(\square)$ by the tensor product structure. With the above norm equivalences, it is easy to see that $\mathfrak{H}^{0,\gamma}(\square)$ is continuously embedded into $\mathfrak{H}^{q,s}(\square)$ if either $q\geq 0$ and $s<\gamma-q$, or if $q<0$ and $s<\gamma$, see also [3]. Furthermore, we can derive the following proposition.
@@ 3.2. Besov Spaces
$|\mathbf{u}|_{\mathbf{b}^{\alpha,p}_{\tau}}\mathrel{\mathrel{\mathop{:}}=}\left[\sum_{m\geq j_{0}}2^{\tau m\alpha}\left[\sum_{\mu\in\hexagon_{m}}|u_{\mu}|^{p}\right]^{\frac{\tau}{p}}\right]^{\frac{1}{\tau}},$ (17)
and $\mathbf{b}^{\alpha,p}_{\tau}$ as the space containing all vectors for which the above quantity (17) is finite. For $\max\{p,\tau\}=\infty$, we apply the usual modifications. Then, the Besov space $B^{\alpha,p}_{\tau}(\square)$ is defined as the space containing all functions $u=\Theta\mathbf{u}$, for which the norm
$\|u\|_{B^{\alpha,p}_{\tau}(\square)}\mathrel{\mathrel{\mathop{:}}=}\|u\|_{L^{p}(\square)}+|\mathbf{u}^{(p)}|_{\mathbf{b}^{\alpha,p}_{\tau}}$
@@
    $\theta_{\mu}\in B^{\beta,p}_{\tau}(\square)$ for some $\beta>\alpha$ and
@@
    $\theta_{\mu}$ has $r>\max\{\alpha,n(1/p-1)\}$ vanishing moments,
@@
As we will see, the classical Besov spaces and also the Besov spaces of dominating mixed smoothness will not be the right spaces for the approximation with respect to a tensor-product basis in an isotropic energy space. For this, we need to consider Besov spaces of hybrid regularity as introduced in [3], which are characterised by the decay of the wavelet coefficients.
@@
$|\mathbf{u}|_{\mathfrak{b}^{q,s,p}_{\tau}}\mathrel{\mathrel{\mathop{:}}=}\left[\sum_{\mathbf{j}\geq\mathbf{j}_{0}}2^{\tau(q|\mathbf{j}|_{\infty}+s|\mathbf{j}|_{1})}\left[\sum_{\boldsymbol{\lambda}\in\boldsymbol{\nabla}_{\mathbf{j}}}\big{|}u_{\boldsymbol{\lambda}}\big{|}^{p}\right]^{\frac{\tau}{p}}\right]^{\frac{1}{\tau}},$ (18)
@@
$\|u\|_{\mathfrak{B}^{q,s,p}_{\tau}(\square)}\mathrel{\mathrel{\mathop{:}}=}\|u\|_{L^{p}(\square)}+|\mathbf{u}^{(p)}|_{\mathfrak{b}^{q,s,p}_{\tau}}$ (19)
@@
The above definitions are given for nonnegative $\alpha$, $q$, and $s$ only, since for $\alpha<0$, $B^{\alpha,p}_{\tau}(\square)$ does not need to be included in $L^{p}(\square)$, and likewise in the hybrid case. Nevertheless, an extension of the seminorms (17) and (18) to negative $\alpha$, $q$, and $s$, respectively, is straightforward.
Remark 3.10.
At the first glance, the expression (18) looks different from that in [3]. This can, however, be explained by the fact that the wavelet bases in [3] are normalised in $L^{\infty}(\square)$. Moreover, if $p=\tau=2$, then we have the identities
@@ 4. Approximation and Interpolation
The aim of this section is to briefly summarise the interpolation between vector spaces and the $N$-term approximation, which is a kind of nonlinear approximation. Our goal is to characterise the approximation spaces for the anisotropic tensor-product wavelet basis. All the results provided here can also be found in [11].
@@ 4.1. Interpolation
First, we want to specify the abstract interpolation between normed vector spaces by means of the $K$-functional. We consider two normed vector spaces, $(V,\|\cdot\|_{V})$ and $(W,\|\cdot\|_{W})$, and we assume that $W$ is continuously embedded in $V$, that is, $\|\cdot\|_{V}\lesssim\|\cdot\|_{W}$. Moreover, let $|\cdot|_{W}$ be a seminorm on $W$. Then, for $t>0$, we define the $K$-functional by
@@
One can show that there holds
$|u|_{(V,W)_{\theta,\tau}}\lesssim\|u\|_{V}^{1-\theta}|u|_{W}^{\theta},\qquad u\in W\subseteq V.$
@@ 4.2. Approximation
Another question is the following: Given a Riesz basis $\Psi=\{\psi_{\lambda}:\lambda\in\nabla\}$ of a space $V$, for which subspace $W\subseteq V$ can we approximate all functions $u\in W$ with at most $N$ terms at the rate $N^{-s}$? This is a topic of nonlinear approximation and, in particular, best $N$-term approximation, as we consider trial spaces
$V_{N}\mathrel{\mathrel{\mathop{:}}=}\left\{\sum_{\lambda\in\Lambda}c_{\lambda}\psi_{\lambda}:c_{\lambda}\in\mathbb{R},\ |\Lambda|\leq N\right\}.$
@@
provided that $\Theta\subseteq B^{q+rn+\varepsilon,\tau}_{\tau}(\square)$ for some $\varepsilon>0$ and that the wavelets admit $\tilde{d}>\max\{q+rn,n(r-\frac{1}{2})\}$ vanishing moments.
@@ 4.3. Approximation with Tensor-Product Wavelets
With the results of the previous two subsections, we can now classify the approximation spaces for any function space $H^{q}(\square)$ with respect to an anisotropic tensor-product wavelet basis.
@@
For $r>0$, $q\in(-\tilde{\gamma},\gamma)$, and $\tau$ such that
$\frac{1}{\tau}=r+\frac{1}{2},$
there holds the Jackson inequality
$\inf_{v_{N}\in V_{N}}\|u-v_{N}\|_{H^{q}(\square)}\lesssim N^{-r}|u|_{\mathfrak{B}^{q,r,\tau}_{\tau}(\square)}.$ (20)
@@
We use an argument similar to [11, 22]. First, we remark that
@@
is a Riesz basis for $H^{q}(\square)$ due to the norm equivalence (12). Thus, the best $N$-term approximation to any function $u$ asymptotically corresponds to the $N$ terms with the largest coefficients in absolute value. Hence, if $\mathbf{u}^{\star}=(\mathbf{u}^{\star}(k))_{k}$ is a descending (in absolute value) reordering of the sequence $(2^{q|\boldsymbol{\lambda}|_{\infty}}u_{\boldsymbol{\lambda}})_{\boldsymbol{\lambda}}$, then there holds
@@
Therefore, it is enough to require that the quantity
$\|\mathbf{u}^{\star}\|_{\ell^{\tau}_{\infty}}\mathrel{\mathrel{\mathop{:}}=}\sup_{k\in\mathbb{N}}k^{\frac{1}{\tau}}|\mathbf{u}^{\star}(k)|$
@@
$\inf_{v_{N}\in V_{N}}\|u-v_{N}\|_{H^{q}(\square)}\lesssim N^{-r}\|\mathbf{u}^{\star}\|_{\ell^{\tau}(\mathbb{N})}.$ (22)
@@
$=|u|_{\mathfrak{B}^{q,r,\tau}_{\tau}(\square)}^{\tau}.$
Together with (22), this implies (20). ∎
Theorem 4.4.
Assume that the assumptions of Theorem 4.3 hold. Then, we also have the Bernstein inequality
$|u_{N}|_{\mathfrak{B}^{q,r,\tau}_{\tau}(\square)}\lesssim N^{r}\|u_{N}\|_{H^{q}(\square)},\qquad u_{N}\in V_{N}.$ (23)
Proof.
We will again make use of the norm equivalence (12) for the wavelet basis. Since $u_{N}\in V_{N}$, we may write $u_{N}$ as a linear combination
@@
$|u_{N}|_{\mathfrak{B}^{q,r,\tau}_{\tau}(\square)}^{\tau}$ $=\sum_{\boldsymbol{\lambda}\in\Lambda}2^{\tau(q|\boldsymbol{\lambda}|_{\infty}+r|\boldsymbol{\lambda}|_{1})}\big{|}u_{\boldsymbol{\lambda}}^{(\tau)}\big{|}^{\tau}=\sum_{\boldsymbol{\lambda}\in\Lambda}2^{\tau(q|\boldsymbol{\lambda}|_{\infty}+r|\boldsymbol{\lambda}|_{1})}\big{|}2^{|\boldsymbol{\lambda}|_{1}\big{(}\frac{1}{2}-\frac{1}{p}\big{)}}u_{\boldsymbol{\lambda}}\big{|}^{\tau}$
@@
With the Jackson and the Bernstein inequality at hand, Theorem 4.1 can be used to characterise the approximation spaces.
Theorem 4.5.
Assume that the assumptions of Theorem 4.3 hold. Then, for $0<s<r$ and $0<\kappa\leq\infty$, the approximation spaces with respect to tensor-product wavelets are given by

### v2

@@ 3.1. Sobolev Spaces
we get a primal basis which is normalised in $L^{p}(\square)$, and a dual basis which is normalised in $L^{p^{\prime}}(\square)$, where $\frac{1}{p}+\frac{1}{p^{\prime}}=1$. With the above notation at hand, every function $u\in L^{2}(\square)$ admits unique expansions
$\begin{aligned} u&=\sum_{\boldsymbol{\lambda}\in\boldsymbol{\nabla}}u_{\boldsymbol{\lambda}}^{(p)}\psi_{\boldsymbol{\lambda}}^{(p)},&\qquad u_{\boldsymbol{\lambda}}^{(p)}&=\big\langle\tilde{\psi}_{\boldsymbol{\lambda}}^{(p)},\,u\big\rangle,\\ u&=\sum_{\mu\in\hexagon}u_{\mu}^{(p)}\theta_{\mu}^{(p)},&\qquad u_{\mu}^{(p)}&=\big\langle\tilde{\theta}_{\mu}^{(p)},u\big\rangle.\end{aligned}$
@@
$\|u\|_{H^{t}(\square)}^{2}\sim\sum_{\boldsymbol{\lambda}\in\boldsymbol{\nabla}}2^{2t|\boldsymbol{\lambda}|_{\infty}}\big|\langle u,\psi_{\boldsymbol{\lambda}}\rangle\big|^{2}\sim\sum_{\mu\in\hexagon}2^{2s|\mu|}\big|\big\langle u,\theta_{\mu}\rangle\big|^{2},\quad-\gamma<t<\tilde{\gamma}$, (15)
cf. [1, 7, 18, 22, 32, 33]. In particular, a function $u$ is contained in an isotropic Sovolev space if the coefficients of its (tensor-product) wavelet expansion decay sufficiently fast. Note that we have slightly abused the notation here as we have omitted the boundary conditions which are imposed by the primal and dual wavelet bases in (14) and (15), respectively. Note that, in accordance with [9], all boundary conditions which do not change within a face of the unit cube can be expressed.
@@
By using the norm equivalences (14) and (15), we see that we can also characterise the spaces $\mathfrak{H}^{q,s}(\square)$ in terms of hyperbolic wavelet coefficients. Note that the following norm equivalences have already been established in [16, 17] for a smaller range of parameters. Those results cover the cases $s\geq 0$ and $0\leq q+s<\gamma$, which is not sufficient to prove the Jackson and Bernstein inequalities in Section 4.3.
@@
$\|u\|_{H^{q+s}([0,1])\otimes H^{s}([0,1])}^{2}$ $\sim\sum_{\boldsymbol{\lambda}\in\boldsymbol{\nabla}}2^{2(q+s)|\lambda_{1}|+2s|\lambda_{2}|}|u_{\boldsymbol{\lambda}}|^{2},$
$\|u\|_{H^{s}([0,1])\otimes H^{q+s}([0,1])}^{2}$ $\sim\sum_{\boldsymbol{\lambda}\in\boldsymbol{\nabla}}2^{2s|\lambda_{1}|+2(q+s)|\lambda_{2}|}|u_{\boldsymbol{\lambda}}|^{2},$
provided that $\tilde{\gamma}<s,q+s<\gamma$. Therfore, as $q+s\geq s$, there holds
$2^{2(q+s)|\lambda_{1}|+2s|\lambda_{2}|}+2^{2s|\lambda_{1}|+2(q+s)|\lambda_{2}|}$ $\sim 2^{2(q+s)|\boldsymbol{\lambda}|_{\infty}+2s\min\{|\lambda_{1}|,|\lambda_{2}|\}}$
@@
where we have used (18) for $q\geq 0$.
@@
Note that we have again used (18) for $q\geq 0$. This shows (17).
@@
We note that the one-sided upper bound in (17) can be extended for $\tilde{d}<s,q+s<\gamma$, whereas the one-sided lower bound can be extended to $\tilde{\gamma}<s,q+s<d$. Herein, $d$ and $\tilde{d}$ denote the order of polynomial exactness of $\Psi$ and $\tilde{\Psi}$, which implies in turn that $\Psi$ and $\tilde{\Psi}$ admit $\tilde{d}$ and $d$ vanishing moments, respectively. Similarly, we can also extend the upper and lower bounds in (18) up to d-d and $\tilde{d}$, respectively.
We note that there holds $\theta_{\mu},\psi_{\boldsymbol{\lambda}}\in\mathfrak{H}^{0,\gamma}(\square)\subseteq\mathfrak{H}^{\gamma,0}(\square)=H^{\gamma}(\square)$ by the tensor-product structure. With the above norm equivalences, it is easy to see that $\mathfrak{H}^{0,\gamma}(\square)$ is continuously embedded into $\mathfrak{H}^{q,s}(\square)$ if either $q\geq 0$ and $s<\gamma-q$, or if $q<0$ and $s<\gamma$, see also [3]. Furthermore, we can derive the following proposition.
@@ 3.2. Besov Spaces
$\|\mathbf{u}\|_{\mathbf{b}^{\alpha,p}_{\tau}}\mathrel{\mathrel{\mathop{:}}=}\left[\sum_{m\geq j_{0}}2^{\tau m\alpha}\left[\sum_{\mu\in\hexagon_{m}}|u_{\mu}|^{p}\right]^{\frac{\tau}{p}}\right]^{\frac{1}{\tau}},$ (19)
and $\mathbf{b}^{\alpha,p}_{\tau}$ as the space containing all vectors for which the above quantity (19) is finite. For $\max\{p,\tau\}=\infty$, we apply the usual modifications. Then, the isotropic Besov space $B^{\alpha,p}_{\tau}(\square)$ is defined as the space containing all functions $u=\Theta\mathbf{u}$, for which the norm
$\|u\|_{B^{\alpha,p}_{\tau}(\square)}\mathrel{\mathrel{\mathop{:}}=}\|\mathbf{u}^{(p)}\|_{\mathbf{b}^{\alpha,p}_{\tau}}$
@@
    $\theta_{\mu}\in B^{\beta,p}_{\tau}(\square)$ for some $\beta>\alpha>\max\{0,n(\nicefrac{{1}}{{p}}-1)\}$ and
@@
    $\theta_{\mu}$ has polynomial order $d>\max\{\alpha,n(\nicefrac{{1}}{{p}}-1)\}$,
@@
As we will see, the classical Besov spaces and also the Besov spaces of dominating mixed smoothness will not be the right spaces for the approximation with respect to a tensor-product basis in an isotropic energy space. For this, we need to consider Besov spaces of hybrid regularity as introduced in [3], which are characterised by the decay of the hyperbolic wavelet coefficients.
@@
$\|\mathbf{u}\|_{\mathfrak{b}^{q,s,p}_{\tau}}\mathrel{\mathrel{\mathop{:}}=}\left[\sum_{\mathbf{j}\geq\mathbf{j}_{0}}2^{\tau(q|\mathbf{j}|_{\infty}+s|\mathbf{j}|_{1})}\left[\sum_{\boldsymbol{\lambda}\in\boldsymbol{\nabla}_{\mathbf{j}}}\big|u_{\boldsymbol{\lambda}}\big|^{p}\right]^{\frac{\tau}{p}}\right]^{\frac{1}{\tau}},$ (20)
@@
$\|u\|_{\mathfrak{B}^{q,s,p}_{\tau}(\square)}\mathrel{\mathrel{\mathop{:}}=}\|\mathbf{u}^{(p)}\|_{\mathfrak{b}^{q,s,p}_{\tau}}$ (21)
@@
At the first glance, the expression (20) looks different from that in [3]. This can, however, be explained by the fact that the wavelet bases in [3] are normalised in $L^{\infty}(\square)$. Moreover, a hyperbolic Littlewood-Paley decomposition of hybrid regularity Besov spaces has recently been provided in [14].
If $p=\tau=2$, then we have the identities
@@ 4. Approximation and Interpolation
The aim of this section is to briefly summarise the interpolation between vector spaces and the best $N$-term approximation, which is a kind of nonlinear approximation. Our goal is to characterise the approximation spaces for the anisotropic tensor-product wavelet basis. All the results provided on interpolation and its interplay with approximation can also be found in [12].
@@ 4.1. Interpolation
First, we want to specify the abstract interpolation between normed vector spaces by means of the $K$-functional, cf. [24]. We consider two normed vector spaces, $(V,\|\cdot\|_{V})$ and $(W,\|\cdot\|_{W})$, and we assume that $W$ is continuously embedded in $V$, that is, $\|\cdot\|_{V}\lesssim\|\cdot\|_{W}$. Moreover, let $|\cdot|_{W}$ be a quasi-seminorm on $W$. Then, for $u\in V$ and $t>0$, we define the $K$-functional by
@@
The term interpolation is justified by the property
$|u|_{(V,W)_{\theta,\tau}}\lesssim\|u\|_{V}^{1-\theta}|u|_{W}^{\theta},\qquad u\in W\subseteq V,$
for example. For a proof and further results on interpolation, consult [36], for instance.
@@ 4.2. Approximation
A question which is closely related to interpolation is the following: Given a Riesz basis $\Psi=\{\psi_{\lambda}:\lambda\in\nabla\}$ of a space $V$, for which subspace $W\subseteq V$ can we approximate all functions $u\in W$ with at most $N$ terms at the rate $N^{-s}$? This is a topic of nonlinear approximation and, in particular, best $N$-term approximation, as we consider trial spaces
$V_{N}\mathrel{\mathrel{\mathop{:}}=}\left\{\sum_{\lambda\in\Lambda}c_{\lambda}\psi_{\lambda}:c_{\lambda}\in\mathbb{C},\ |\Lambda|\leq N\right\}.$
@@
provided that $\Theta\subseteq B^{q+rn+\varepsilon,\tau}_{\tau}(\square)$ for some $\varepsilon>0$ and that the wavelets admit polynomial exactness $d>\max\{q+rn,n(r-\frac{1}{2})\}$.
@@ 4.3. Approximation with Tensor-Product Wavelets
With the results of the previous two subsections, we can now classify the approximation spaces for any function space $H^{q}(\square)$ with respect to an anisotropic tensor-product wavelet basis. Note that the following proof is implicitly contained in [20, 21], but simplifies a lot in the current setting.
@@
Assume that $r>0$, $\gamma<q<\gamma$, and
$\frac{1}{\tau}=r+\frac{1}{2}.$
Then, for $0<s<r$ and $0<\kappa\leq\infty$, the approximation spaces with respect to tensor-product wavelets are given by
@@
In accordance with Theorem 4.1, we need to show the Jackson and Bernstein inequalities
$\inf_{v_{N}\in V_{N}}\|u-v_{N}\|_{H^{q}(\square)}$ $\lesssim N^{-r}\|u\|_{\mathfrak{B}^{q,r,\tau}_{\tau}(\square)},$ $u$ $\in\mathfrak{B}^{q,r,\tau}_{\tau}(\square),$ (23)
$\|u_{N}\|_{\mathfrak{B}^{q,r,\tau}_{\tau}(\square)}$ $\lesssim N^{r}\|u_{N}\|_{H^{q}(\square)},$ $u_{N}$ $\in V_{N},$ (24)
to conclude the statement of the theorem. In order to do achieve this, we use an argument similar to [12, 29].
Let us show the Jackson inequality (23) first. We note that
@@
is a Riesz basis for $H^{q}(\square)$ due to the norm equivalence (14). Thus, the best $N$-term approximation to any function $u$ asymptotically corresponds to the $N$ terms with the largest coefficients in absolute value. Hence, if $\mathbf{u}^{\star}=(\mathbf{u}^{\star}(k))_{k}$ is a descending (in absolute value) reordering of the sequence $(2^{q|\boldsymbol{\lambda}|_{\infty}}u_{\boldsymbol{\lambda}})_{\boldsymbol{\lambda}}$, there holds
@@
Therefore, it is equivalent to require that the quantity
$\|\mathbf{u}^{\star}\|_{\ell^{\tau}_{\infty}}\mathrel{\mathrel{\mathop{:}}=}\sup_{k\in\mathbb{N}}k^{\frac{1}{\tau}}|\mathbf{u}^{\star}(k)|,\qquad\frac{1}{\tau}=r+\frac{1}{2},$
@@
$\inf_{v_{N}\in V_{N}}\|u-v_{N}\|_{H^{q}(\square)}^{2}$ $\lesssim\sum_{k=N+1}^{{}^{\infty}}|\mathbf{u}^{\star}(k)|^{2}\lesssim\|\mathbf{u}^{\star}\|_{\ell^{\tau}_{\infty}(\mathbb{N})}^{2}\int_{N}^{\infty}t^{-\frac{2}{\tau}}\operatorname{d}\!t$
$\lesssim N^{-2r}\|\mathbf{u}^{\star}\|_{\ell^{\tau}(\mathbb{N})}^{2}.$ (26)
@@
$=\|u\|_{\mathfrak{B}^{q,r,\tau}_{\tau}(\square)}^{\tau}.$
Together with (26), this implies (23).
Also for the Bernstein inequality (24), we will make use of the norm equivalence (14). Since $u_{N}\in V_{N}$, we may write $u_{N}$ as a linear combination
@@
$\|u_{N}\|_{\mathfrak{B}^{q,r,\tau}_{\tau}(\square)}^{\tau}$ $=\sum_{\boldsymbol{\lambda}\in\Lambda}2^{\tau(q|\boldsymbol{\lambda}|_{\infty}+r|\boldsymbol{\lambda}|_{1})}\big|u_{\boldsymbol{\lambda}}^{(\tau)}\big|^{\tau}=\sum_{\boldsymbol{\lambda}\in\Lambda}2^{\tau(q|\boldsymbol{\lambda}|_{\infty}+r|\boldsymbol{\lambda}|_{1})}\big|2^{|\boldsymbol{\lambda}|_{1}\big(\frac{1}{2}-\frac{1}{p}\big)}u_{\boldsymbol{\lambda}}\big|^{\tau}$